# RadZero: Similarity-Based Cross-Attention for Explainable Vision-Language Alignment in Radiology with Zero-Shot Multi-Task Capability

### v1

@@ Abstract
Recent advancements in multi-modal models have significantly improved vision-language alignment in radiology. However, existing approaches struggle to effectively utilize complex radiology reports for learning, rely on low-resolution images, and offer limited interpretability in attention mechanisms. To address these challenges, we introduce RadZero, a novel similarity-based cross-attention framework for vision-language alignment in radiology with zero-shot multi-task capability. RadZero leverages large language models to extract minimal semantic sentences from radiology reports and employs a multi-positive contrastive learning strategy to effectively capture relationships between images and multiple relevant textual descriptions. It also utilizes a pre-trained vision encoder with additional trainable Transformer layers, allowing efficient high-resolution image processing. By computing similarity between text embeddings and local image patch features, RadZero enables zero-shot inference with similarity probability for classification and pixel-level cross-modal similarity maps for grounding and segmentation. Experimental results on public chest radiograph benchmarks show that RadZero outperforms state-of-the-art methods in zero-shot classification, grounding, and segmentation. Furthermore, cross-modal similarity map analysis highlights its potential for improving explainability in vision-language alignment. Additionally, qualitative evaluation demonstrates RadZero’s capability for open-vocabulary semantic segmentation, further validating its effectiveness in medical imaging.
@@ 1 Introduction
Figure 1: Summary of RadZero’s capabilities. (a) Zero-shot multi-task performance (b) Open-vocabulary semantic segmentation. The original CXR images and the meanings of the abbreviations can be found in the Appendix.
@@
Recent advances in deep learning have significantly transformed medical imaging, leading to numerous studies proposing computer-aided diagnosis [15, 8, 23]. However, obtaining high-quality manual annotations for medical images remains a major challenge. In the natural image domain, the emergence of vision-language models [27, 37, 36] has significantly reduced the reliance on exhaustive manual labels. These models learn the relationships between language and vision using only image-text pairs without label supervision, demonstrating strong performance even on zero-shot tasks such as classification and retrieval. Following these advancements, vision-language modeling has been rapidly progressing in the field of medical imaging, including chest X-rays (CXRs). Several studies focused on learning effective representations [34, 39, 40], demonstrating zero-shot capabilities [19, 22] without the need for task-specific annotations.
Despite these promising advances, medical vision-language alignment still faces several challenges. Effectively utilizing complex radiology reports for learning remains difficult. Previous attempts to address the issue include local alignment at the word or token level [12, 1], extracting clinical entities [39, 34], and LLM-based prompt alignment [19]. However, these approaches still suffer from limitations, such as failing to segment text embeddings into appropriate semantic units or the need for random selection during training, which leads to inefficient learning. Additionally, prior studies focusing on zero-shot tasks [19, 34] have relied on low-resolution images, which degrade performance on fine-grained tasks requiring precise localization. An even more critical challenge is ensuring explainability, as medical imaging models require interpretable outputs for clinical use. While attention maps [34, 19, 39] and dot-product similarities [12] are commonly used explainable features, they are not suitable for image-text similarity measures, limiting their effectiveness in interpretability.
@@
To address these challenges, we propose RadZero, a novel vision-language alignment framework in radiology with zero-shot multi-task capability. We employ multi-positive contrastive learning [20] to effectively leverage multiple sentences associated with each image in an image-report pair. To use high-resolution images for improved performance on fine-grained zero-shot tasks, we adopt the idea of freezing a pretrained image encoder [36] and adding trainable Transformer layers for efficient training [18]. The key innovation of RadZero is the similarity-based cross-attention mechanism, which directly computes the similarity between text descriptions and local image patches. Cross-modal similarity map analysis demonstrates that these maps significantly enhance explainability by providing clear and interpretable visual reasoning for the model’s decisions. Additionally, we demonstrate the potential for open-vocabulary semantic segmentation, which can be achieved by simply applying a threshold to the similarity map. Experimental results on public chest radiograph benchmarks show that RadZero outperforms state-of-the-art (SOTA) models in zero-shot classification, grounding, and segmentation. Figure 1 summarizes the capabilities demonstrated by RadZero.
@@ 2.1 General vision-language alignment
Contrastive learning for vision–language alignment with large-scale image–text pairs has been actively studied. CLIP [27] demonstrated that this approach enables strong zero-shot classification by directly aligning images and text. LiT [36] proposed freezing the pretrained vision encoder during contrastive training, preserving fine-grained visual features and further improving zero-shot performance. dino.txt [18] extended this paradigm by incorporating additional Transformer [31] layers on top of a pretrained DINOv2 [25], training only a lightweight module while keeping the vision encoder frozen. Additionally, it fused global and patch-averaged embeddings, enabling patch-level similarity computation with text and facilitating open-vocabulary semantic segmentation. UniCLIP [20] introduced a multi-positive NCE (MP-NCE) loss, which independently computes the contribution of multiple positive pairs per image. Building on these advances, our approach integrates a frozen, fine-grained vision encoder with trainable Transformer layers, following strategies from LiT and dino.txt. Additionally, we leverage MP-NCE loss to effectively align images with multiple text representations.
Figure 2: The overall framework of RadZero. (a) Finding-sentence extraction using an LLM. (b) Computation of the similarity logit, $l_{ij}^{n}$, between image $I_{i}$ and finding-sentence $S_{j}^{n}$. W-sum and cos-sim denote weighted sum and cosine similarity, respectively. (c) Computation of MP-NCE loss ($\mathcal{L}_{I}$) and InfoNCE loss ($\mathcal{L}_{T}$) from the similarity logit matrix.
@@ 2.2 Vision-language alignment in radiology
ConVIRT [40] initially applied contrastive learning to align CXR images with radiology reports. GLoRIA [12] focused on local alignment by introducing cross-attention between word-level text embeddings and patch-level image features, whereas MGCA [32] proposed a multi-granularity approach to capture relationships at the disease, instance, and pathological region levels. BioViL [3] refined the language model architecture to better handle radiology reports and BioViL-T [1] extended further by incorporating prior images during training. MedKLIP [34] utilized RadGraph [14] to extract triplets from reports and incorporated them into the training process, achieving notable performance in both zero-shot and fine-tuning settings. Similarly, KAD [39] employed RadGraph for entity extraction and used contrastive loss along with disease-specific queries in a cross-attention framework. More recently, CARZero [19] introduced an LLM-based prompt alignment to standardize diagnostic expressions in radiology reports, leveraging cross-attention alignment to achieve reliable zero-shot classification and grounding performance.
2.3 Fine-grained similarity for explainability
Although explainability is essential in medical vision-language alignment, the explainable features for image-text similarity proposed by previous studies have been insufficient. BioViL [3] computed patch-level similarity but without cross-attention on visual patches, limiting spatial alignment. GLoRIA [12] and BioViL-T [1] relied on word- or token-level embeddings, which lack semantic richness. Attention maps, though selected as explainable features by numerous zero-shot studies [39, 34, 19], do not constitute a valid measure of similarity between image and text. Due to the softmax activation, even unrelated image–text pairs yield high values at certain points in the map. The simple removal of softmax to use the raw logits is not a desirable solution, as the logits are neither scaled nor centered at zero. Moreover, the vector norms of query and key embeddings vary according to the image or text, causing fluctuations in the similarity value scale across different image–text pairs, an issue that dot-product similarity-based approaches [12] also encounter.
In contrast, our approach significantly improves explainability by employing similarity-based cross-attention between visual patches and text embeddings extracted from minimal semantic units via an LLM. The resulting feature maps are clear measures of fine-grained text–image similarity, providing intuitive and consistent explanations.
@@ 3.1 Finding-sentence extraction
Radiology reports contain a mixture of various types of information, including clinical history, observations, comparative analysis with prior studies, and diagnostic impressions. When using image and full report pairs for training, it becomes challenging to encode the complex radiology report into a single text embedding. CARZero [19] addressed this issue by leveraging an LLM to extract sentences from radiology reports and proposed a prompt alignment strategy using the template “There is [disease]” to ensure consistency between training and inference. Similarly, we utilize an LLM to extract such sentences, which we refer to as finding-sentences. We design a prompt that includes the positional information of the findings within a predefined template, such as “There is [finding] of [location]”. Each finding-sentence is segmented into the minimal semantic unit that contains the name of the finding, its presence (or uncertainty), and location information. The prompt used for finding-sentence extraction is in Appendix D. For each image, multiple finding-sentences are matched and used during training. An example of the finding-sentence extraction is illustrated in Figure 2 (a). For zero-shot inference, we apply prompt alignment by prepending “There is” to text descriptions such as findings and anatomical regions.
@@ 3.2.1 Model architecture
To leverage the advantages of vision encoder pre-training, we adopt the approach of LiT [36] by freezing a pre-trained vision encoder in contrastive learning. In Vision Transformers [9] such as DINOv2 [25], interpolating the positional embeddings allows for increased input image resolution [29]. Building on this property, we train our model with high-resolution images. To embed the output of the vision encoder, we add trainable Transformer layers, as proposed by Jose et al. [18]. For the text encoder, we use a pre-trained Sentence-BERT [28], which is fine-tuned during training, to extract embeddings for each finding-sentence. The model architecture is illustrated in Figure 2 (b).
Figure 3: Numpy-like pseudocode implementation of similarity-based cross-attention.
3.2.2 Similarity-based cross-attention
To address the issues discussed in Sec. 2.3, we propose a cosine similarity-based cross-attention mechanism for computing similarity logits. By directly employing cosine similarity between the text and visual patch embeddings, we obtain cross-modal similarity scores that are well-defined in range and centered at zero. This consistent scaling allows for fair comparisons across different image–text pairs and significantly enhances explainability through the visualization of similarity maps. It also enables single thresholding, new possibilities for open-vocabulary semantic segmentation.
Specifically, let the size of a mini-batch be $B$. The $i$-th image $I_{i}$ ($i\in\{1,\dots,B\}$) is paired with $N_{i}$ finding-sentences. For $I_{i}$ and the $n$-th finding-sentence of the $j$-th image $I_{j}$ ($n\in\{1,\dots,N_{j}\}$), denoted by $S_{j}^{n}$, the algorithm to compute the patch-level similarity map $M_{ij}^{n}$ and the similarity logit $l_{ij}^{n}$ is summarized in Figure 3 as a numpy-like pseudocode.
From the vision encoder, we obtain a vision token sequence $V_{i}$, composed of the [CLS] token embedding and visual patch embeddings. The subsequent cross-attention operation takes $T_{j}^{n}$, the text embedding of $S_{j}^{n}$, as the query and $V_{i}$ as both the key and value. We compute the cosine similarity between $T_{j}^{n}$ and each token in $V_{i}$, and to ensure distinguishability, we multiply it by a temperature parameter. The softmax activation transforms these scaled similarity scores into attention probabilities. The weighted sum of the attention probabilities and the vision tokens yields a refined vision embedding. Finally, we compute its cosine similarity with $T_{j}^{n}$, multiply it by a temperature parameter, and obtain the overall similarity logit $l_{ij}^{n}$. Additionally, we obtain the patch-level cross-modal similarity map $M_{ij}^{n}$ by excluding the first element (corresponding to the [CLS] token) from the similarity scores.
3.3 Training objectives
@@
$\mathcal{L}_{I}=-\frac{1}{N_{T}}\sum_{i=1}^{B}\sum_{n=1}^{N_{i}}\log\frac{e^{l_{ii}^{n}}}{e^{l_{ii}^{n}}+\sum\limits_{j\neq i}^{B}\sum\limits_{m=1}^{N_{j}}e^{l_{ij}^{m}}}$ (1)
@@
$\mathcal{L}_{T}=-\frac{1}{N_{T}}\sum_{i=1}^{B}\sum_{n=1}^{N_{i}}\log\frac{e^{l_{ii}^{n}}}{e^{l_{ii}^{n}}+\sum\limits_{j\neq i}^{B}e^{l_{ji}^{n}}}$ (2)
The final objective function is the sum of $\mathcal{L}_{I}$ and $\mathcal{L}_{T}$:
$\mathcal{L}=\mathcal{L}_{I}+\mathcal{L}_{T}.$ (3)
Figure 4: Zero-shot inference pipeline of RadZero.
@@ 3.4 Zero-shot inference
The similarity logit between an image $I_{i}$ and a text sentence $S_{j}^{n}$, denoted by $l_{ij}^{n}$, is converted into the final similarity probability $\hat{l}_{ij}^{n}$ by applying a sigmoid function, which is then used for classification. For zero-shot grounding and segmentation, pixel-level probability values are required. Since the patch-level cross-modal similarity map $M_{ij}^{n}$ is computed after image pre-processing such as padding and resizing, we restore it to the original size using linear interpolation while accounting for these transformations. Then, a sigmoid function is applied to obtain the pixel-level cross-modal similarity map ${\hat{M}}_{ij}^{n}$, which we simply refer to as the similarity map. The zero-shot inference process is illustrated in Figure 4.
3.5 Explainability
In the cross-attention and logit computation process, we do not modify the embedding space beyond applying $L_{2}$ normalization and adjusting the temperature. Since similarity is computed using cosine similarity between vision patches and text embeddings, the final similarity map ${\hat{M}}_{ij}^{n}$ can be directly interpreted as the similarity between each image pixel and the text, significantly improving the model’s explainability. For instance, when the similarity between a text prompt and an image is low, it directly reflects weak alignment between the vision token sequence and the text embedding, as the final similarity logit is derived from a weighted sum of vision tokens. Conversely, when the similarity is high, it indicates that the text embedding closely aligns with either the entire vision token sequence or specific visual patches. This approach allows for a transparent visualization of how similarity logits are computed through the similarity map. In the context of disease diagnosis, the model can explicitly reveal how conclusions are derived, greatly enhancing its explainability and making its decision-making process more interpretable.

### v2

@@ Abstract
Recent advancements in multi-modal models have significantly improved vision-language (VL) alignment in radiology. However, existing approaches struggle to effectively utilize complex radiology reports for learning and offer limited interpretability through attention probability visualizations. To address these challenges, we introduce RadZero, a novel framework for VL alignment in radiology with zero-shot multi-task capability. A key component of our approach is VL-CABS (Vision-Language Cross-Attention Based on Similarity), which aligns text embeddings with local image features for interpretable, fine-grained VL reasoning. RadZero leverages large language models to extract concise semantic sentences from radiology reports and employs multi-positive contrastive training to effectively capture relationships between images and multiple relevant textual descriptions. It uses a pre-trained vision encoder with additional trainable Transformer layers, allowing efficient high-resolution image processing. By computing similarity between text embeddings and local image patch features, VL-CABS enables zero-shot inference with similarity probability for classification, and pixel-level VL similarity maps for grounding and segmentation. Experimental results on public chest radiograph benchmarks show that RadZero outperforms state-of-the-art methods in zero-shot classification, grounding, and segmentation. Furthermore, VL similarity map analysis highlights the potential of VL-CABS for improving explainability in VL alignment. Additionally, qualitative evaluation demonstrates RadZero’s capability for open-vocabulary semantic segmentation, further validating its effectiveness in medical imaging.
@@ 1 Introduction
Recent advancements in deep learning have significantly impacted medical imaging, leading to numerous studies on computer-aided diagnosis [15, 8, 24, 35]. However, acquiring high-quality manual annotations remains a key challenge. In contrast, vision-language (VL) models (VLMs) in the natural image domain [28, 39, 38] have reduced reliance on manual labeling by learning from image-text pairs without explicit supervision, achieving strong zero-shot performance in tasks like classification and retrieval. Building on this progress, VLMs have been increasingly explored in medical imaging, including chest X-rays (CXRs). Several studies have demonstrated effective representation learning [36, 41, 42] and zero-shot capabilities [19, 23] without task-specific annotations.
Figure 1: Comparison of attention maps and the proposed VL similarity map for visualizing VL alignment. (a) While traditional attention maps inevitably exhibit high values at certain points due to softmax activation, the proposed VL similarity maps yield low values for unrelated image-text pair. (b) Their fixed scale, originating from cosine similarity, enables open-vocabulary semantic segmentation through simple thresholding.
Despite these advance, current medical VLMs underutilize the rich semantics of radiology image–report pairs. Prior methods—such as word-level alignment [12, 1], clinical entity extraction [41, 36], and using LLM prompts [19]—face limitations, struggle with poor text embedding segmentation and inefficient training due to random sampling. An effective solution should 1) decompose reports into semantically minimal, clinically meaningful sentences, 2) embed each sentence independently, and 3) learn from multiple sentence–image pairs per study to to fully leverage supervision.
Moreover, reliable explainability is critical for the clinical adoption of medical VLMs. Attention maps are used as explainable features by most of the recent research [36, 19, 41]. However, while an attention map (Figure 1(a), top) can indicate where the model is focusing, it does not provide an explanation for why it is attending to those regions. This limitation can be addressed by computing pixel-level image–text similarity, which enables more fine-grained and transparent explanations.
Figure 2: Zero-shot multi-task performance. Each score is averaged over multiple datasets per task.
@@
To overcome these shortcomings, we propose RadZero, a novel VL alignment framework for radiology with zero-shot multi-task capabilities. RadZero employs multi-positive contrastive learning [20] to incorporate multiple sentences per image-report pair. RadZero’s core innovation, Vision-Language Cross-Attention Based on Similarity (VL-CABS), directly computes cosine similarity between text descriptions and local image patches. Unlike traditional attention maps, the resulting VL similarity maps offer clearer visual reasoning by maintaining low values for unrelated image-text pairs (Figure 1 (a), bottom). This enhances interpretability and enables open-vocabulary semantic segmentation via simple thresholding. RadZero also supports high-resolution inputs by freezing a pre-trained image encoder [38] and adding trainable Transformer layers [18], further boosting performance on fine-grained zero-shot tasks. Experiments on public chest radiograph benchmarks demonstrate that RadZero outperforms state-of-the-art (SOTA) models in various zero-shot tasks (Figure 2), while qualitative analyses reveal its enhanced explainability and potential for open-vocabulary semantic segmentation.
@@ 2.1 General vision-language alignment
Contrastive learning for vision–language alignment with large-scale image–text pairs has been actively studied. CLIP [28] demonstrated that directly aligning images and text enables strong zero-shot classification. LiT [38] proposed freezing the pre-trained vision encoder during contrastive training, preserving fine-grained visual features and further enhancing zero-shot performance. dino.txt [18] extended this framework by adding Transformer [32] layers on top of a pre-trained DINOv2 [26], training only a lightweight module while keeping the vision encoder frozen. Additionally, it fused global and patch-averaged embeddings, enabling patch-level similarity computation with text and supporting open-vocabulary semantic segmentation. UniCLIP [20] introduced a multi-positive NCE (MP-NCE) loss, which independently evaluates multiple positive pairs per image. Building on these advances, our approach integrates a frozen, fine-grained vision encoder with trainable Transformer layers, following LiT and dino.txt. We also adopt MP-NCE loss to align images with multiple text representations effectively.
@@ 2.2 Vision-language alignment in radiology
Since the introduction of contrastive learning in radiology [42], aligning CXR images with radiology reports has become an active research area. GLoRIA [12] focused on local alignment using cross-attention between word-level text embeddings and patch-level image features. MGCA [33] employed both report-level and token-level embeddings to extract multi-granular features, and BioViL-T [1] similarly relied on token-level embeddings. Nevertheless, segmenting reports into individual words or tokens often fails to capture their full semantic meaning.
Due to the complexity of medical image–report relationships, alignment interpretability is essential for clinical use and is commonly addressed using attention maps. MedKLIP [36] and KAD [41], for example, used RadGraph [14] to extract report features and employed attention maps for tasks like grounding and segmentation. In addition to VL alignment, G2D [21] G2D [21] aggregates attention maps in addition to VL alignment to generate pseudo masks, which are used as pixel-wise pretext supervisory signals during pre-training. CARZero [19] also used attention maps when leveraging cross-attention alignment for zero-shot tasks, incorporating LLM-based prompt alignment to standardize reports. Despite their utility, attention maps have limitations: they often highlight irrelevant regions due to softmax activation, but removing softmax is not ideal as raw logits are unnormalized and uncentered. Additionally, variation in the norms of query and key embeddings leads to inconsistent similarity values across different image-text pairs. An example and a detailed discussion of the limited explainability of attention maps are provided in Appendix A. In contrast, our approach enhances explainability with VL-CABS, aligning visual patches and text embeddings. The resulting maps offer intuitive and consistent measures of fine-grained image-text similarity.
@@ 3.1 Finding-sentence extraction
Radiology reports contain diverse types of information, including clinical history, observations, comparative analysis with prior studies, and diagnostic impressions. Encoding the entire report into a single text embedding often fails to capture this complexity. CARZero [19] addressed this by using an LLM to extract relevant sentences and introducing a prompt alignment strategy based on the template “There is [disease]” for consistency between training and inference. Similarly, we use an LLM to extract such sentences, which we refer to as finding-sentences. These are generated using a prompt that follows a predefined structure, such as “There is [finding] of [location],” and are segmented into minimal semantic units containing the finding name, presence (or uncertainty), and location. The full prompt is provided in Appendix D. Each image is paired with multiple finding-sentences during training, as illustrated in Figure 3 (a). For zero-shot inference, we apply prompt alignment by prepending “There is” to text descriptions of findings and anatomical regions.
Figure 3: The overall framework of RadZero. (a) Finding-sentence extraction using an LLM. (b) Computation of the similarity logit, $l_{ij}^{n}$, between image $I_{i}$ and finding-sentence $S_{j}^{n}$. W-sum and cos-sim denote weighted sum and cosine similarity, respectively. (c) Computation of MP-NCE loss ($\mathcal{L}_{I}$) and InfoNCE loss ($\mathcal{L}_{T}$) from the similarity logit matrix. (d) Zero-shot inference pipeline.
@@ 3.2.1 Model architecture
To leverage the advantages of vision encoder pre-training, we adopt the approach of LiT [38] by freezing a pre-trained vision encoder in contrastive learning. In Vision Transformers [9] such as DINOv2 [26], interpolating the positional embeddings allows for increased input image resolution [30]. Building on this property, we train our model with high-resolution images. To embed the output of the vision encoder, we add $K$ trainable Transformer layers, as proposed by Jose et al. [18]. For the text encoder, we use a pre-trained Sentence-BERT [29], which is fine-tuned during training, to extract embeddings for each finding-sentence. The model architecture is illustrated in Figure 3 (b).
3.2.2 Vision-language cross-attention based on similarity
We propose VL-CABS, a cosine similarity-based cross-attention mechanism for computing similarity logits. By directly employing cosine similarity between the text and visual patch embeddings, we obtain VL similarity scores that are well-defined in range and centered at zero. This consistent scaling allows for fair comparisons across different image–text pairs and significantly enhances explainability through the visualization of VL similarity maps. It also enables single thresholding, offering new possibilities for open-vocabulary semantic segmentation.
The proposed method operates on a mini-batch of size $B$, pairing each image $I_{i}$ ($i=1,\dots,B$) with $N_{i}$ associated finding-sentences $\{S_{i}^{n}\}_{n=1}^{N_{i}}$. Each image is processed by a vision encoder $f_{v}$ followed by trainable layers $f_{a}$, yielding a sequence of embeddings $[\,v_{i0},\,v_{i1},\dots,v_{iL}\,]=f_{a}(f_{v}(I_{i})),$ where $v_{i0}\in\mathbb{R}^{D}$ corresponds to the [CLS] token and $\{v_{ik}\}_{k=1}^{L}\subset\mathbb{R}^{D}$ are patch embeddings, with $D$ denoting the embedding dimension and $L$ the total number of patches. Each finding-sentence $S_{j}^{n}$ is encoded into a sentence-level embedding $t_{j}^{n}=f_{t}(S_{j}^{n})\in\mathbb{R}^{D}$ using a text encoder $f_{t}$.
To compute VL similarity, we ℓ2-normalize all embeddings as $\bar{v}_{ik}=v_{ik}/\|v_{ik}\|_{2}$ and $\bar{t}_{j}^{n}=t_{j}^{n}/\|t_{j}^{n}\|_{2}$, and calculate the scaled cosine similarity between each patch and sentence as $s_{ijk}^{n}=(\bar{v}_{ik})^{\top}\bar{t}_{j}^{n}\cdot\exp(\tau)\quad(k=0,\dots,L),$ where $\tau$ is a learnable temperature. These scores are converted into attention weights using softmax over the patch index $k$: $a_{ijk}^{n}=\exp(s_{ijk}^{n})/{\sum_{m=0}^{L}\exp(s_{ijm}^{n})}.$ A sentence-specific attended vision embedding is computed as the weighted sum $v_{ij}^{n}=\sum_{k=0}^{L}a_{ijk}^{n}v_{ik}$, which is then ℓ2-normalized as $\bar{v}_{ij}^{n}=v_{ij}^{n}/\|v_{ij}^{n}\|_{2}$. The global similarity logit between image $I_{i}$ and sentence $S_{j}^{n}$ is given by $l_{ij}^{n}=({\bar{v}_{ij}^{n}})^{\top}\bar{t}_{j}^{n}\cdot\exp(\tau).$ The corresponding patch-level similarity map is $M_{ij}^{n}=[\,s_{ij1}^{n},\dots,s_{ijL}^{n}\,]$.
3.3 Multi-positive contrastive learning
@@
$\mathcal{L}_{I}=-\frac{1}{N_{T}}\sum_{i=1}^{B}\sum_{n=1}^{N_{i}}\log\frac{\exp(l_{ii}^{n})}{\exp(l_{ii}^{n})+\sum\limits_{j\neq i}^{B}\sum\limits_{m=1}^{N_{j}}\exp(l_{ij}^{m})}$ (1)
@@
$\mathcal{L}_{T}=-\frac{1}{N_{T}}\sum_{i=1}^{B}\sum_{n=1}^{N_{i}}\log\frac{\exp(l_{ii}^{n})}{\exp(l_{ii}^{n})+\sum\limits_{j\neq i}^{B}\exp(l_{j,i}^{n})}$ (2)
The final objective function is the sum of $\mathcal{L}_{I}$ and $\mathcal{L}_{T}$: $\mathcal{L}=\mathcal{L}_{I}+\mathcal{L}_{T}.$
@@ 3.4 Zero-shot inference
The similarity logit between an image $I_{i}$ and a sentence $S_{j}^{n}$, denoted by $l_{ij}^{n}$, is converted into a similarity probability $\hat{l}_{ij}^{n}=\sigma(l_{ij}^{n})$ via a sigmoid function, and used for zero-shot classification.
For grounding and segmentation, we reshape the patch-level similarity map $M_{ij}^{n}=[s_{ij1}^{n},\dots,s_{ijL}^{n}]$ into a $\sqrt{L}\times\sqrt{L}$ square map, and resize it to the original image resolution via bilinear interpolation. To account for preprocessing such as padding and resizing, this interpolation is applied accordingly. A final element-wise sigmoid activation is applied to obtain the pixel-level similarity map $\hat{M}_{ij}^{n}=\sigma(\text{bilinear}(M_{ij}^{n}))$, which we refer to as the VL similarity map, used for zero-shot grounding and segmentation. The VL similarity map is derived from the cosine similarity between vision patches and text embeddings, and since we do not modify the embedding space beyond applying ℓ2-normalization and adjusting the temperature, it can be directly interpreted as the similarity between each image pixel and the text. The zero-shot inference process is illustrated in Figure 3 (d).